# Towards Federated RLHF with Aggregated Client Preference for LLMs

### v2

@@ 1 Introduction
Large language models (LLMs) exhibit broad knowledge coverage as they are pretrained on a large corpus. To elevate the quality of their generated content to match the professionalism and readability of human writing, a common approach is to fine-tune these models using reinforcement learning with human feedback (RLHF) Ziegler et al. (2019); Christiano et al. (2017); Ouyang et al. (2022). This process relies on preference datasets, which are constructed through two main methods: human evaluation Bai et al. (2022); Ganguli et al. (2022); Stiennon et al. (2020) and ChatGPT-based ranking Dubois et al. (2024b). In the human-effort approach, a set of instructions (a.k.a. prompts or user queries) is paired with multiple model completions (a.k.a. generated responses), and a team of labelers ranks completions of each instruction from best to worst. In the ChatGPT-based method, the dataset is developed by feeding ChatGPT with multiple inputs encompassing an instruction and a pair of completions, where it selects the superior completion for each input. As LLMs are deployed to serve diverse users, there could be a gap between the preferences of real-world users and those of lablers/ChatGPT, hindering the LLM’s ability to generate responses that align with the users’ preferences.
@@
Therefore, there is a need for a preference dataset that accurately reflects real world users’ preferences in order to enhance the ability of LLM in content generation.
@@
To address privacy concerns, we propose utilizing Federated Learning (FL) techniques to enable large-scale preference collection from diverse real-world users without requiring them to transmit their preference data to a central server. In our design, we adopt FedAvg, a well-established FL algorithm Konečnỳ et al. (2016); McMahan et al. (2017), to learn user preferences without directly collecting their data. In line with the reward model training used in RLHF Stiennon et al. (2020); Ouyang et al. (2022); Rafailov et al. (2023), each LLM user works as a FL client and locally trains a reward model that can evaluate the quality of model completions. The server then aggregates these client-trained reward models into a global model. While this approach appears effective, we observe the following two limitations during training:
  • Excessive Computation Overhead: Typically, the reward model is a regression model designed to output a scalar value that represents the quality of a model completion Stiennon et al. (2020). Its optimization is based on comparing reward differences between preferred and dispreferred completions—preferred completions should receive higher rewards. However, during optimization, each data sample requires the retention of two computation graphs (one for the preferred and one for the dispreferred completion) throughout the forward and backward passes. This results in significant computational overhead and heavy GPU demands.
@@ 2 Related Work
Recent studies have increasingly focused on fine-tuning large language models (LLMs) using federated datasets Sun et al. (2024); Ye et al. (2024); Zhang et al. (2023a); Yi et al. (2023); Zhang et al. (2023b). However, these approaches often suffer from high computation and communication costs due to the necessity of training and synchronizing the model with clients. To mitigate these issues, lightweight methods such as black-box fine-tuning Sun et al. (2023); Lin et al. (2023) and offsite-tuning Wu et al. (2024a); Kuang et al. (2023) have emerged. Despite their advancements, these methods primarily focus on fine-tuning LLMs for specific downstream tasks, neglecting user preferences in the generated responses. A recent benchmark, OpenFedLLM Ye et al. (2024), introduces FedDPO, which allows federated clients to optimize their local LLMs using DPO loss. While this approach can potentially align LLMs with human preferences, it faces three key challenges: excessive computational overhead, preference heterogeneity, and the risk of reward hacking. To address these limitations, our work aims to enable LLMs alignment with a feasible and sustainable training framework in FL.
@@
RLHF typically involves supervised fine-tuning, reward modeling, and reward optimization, initially proposed by Christiano et al. (2017). Proximal Policy Optimization (PPO) Schulman et al. (2017) is a common RLHF algorithm, yet it struggles with instability, inefficiency, and high resource demands Choshen et al. (2019); Engstrom et al. (2020). These challenges have led to the development of alternative methods, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and others Dong et al. (2023); Zhao et al. (2023); Azar et al. (2024); Ethayarajh et al. (2024); Gulcehre et al. (2023), which offer more stable and efficient solutions. However, these methods typically operate within a centralized training framework, where the LLM owner retains control over the preference data. In contrast, our work aims to expand data sources and integrate real user preferences without directly collecting their personal data.
@@ 3.1 Federated Learning (FL)
FL is a distributed training paradigm where a server coordinates various clients toward the same goal, i.e., training a generalized model $\phi\in\mathbb{R}^{d}$ Konečnỳ et al. (2016); McMahan et al. (2017). Consider an FL system with $M$ clients. Denote the weight of client $m$ as $p_{m}$ such that $\sum_{m\in[M]}p_{m}=1$, and we aim to optimize the following objectives:
$\min_{\phi\in\mathbb{R}^{d}}\quad F(\mathbf{\phi})\overset{\triangle}{=}\sum_{m\in[M]}p_{m}F_{m}(\mathbf{\phi})$ (1)
where $F_{m}(\phi)$ is the expected loss on client $m$ given the model $\phi$. As a classical FL algorithm, FedAvg can solve the optimization problem by multiple communications between the server and the clients, followed by a number of recent works Wang et al. (2022); He et al. (2023); Fang et al. (2024); Wang et al. (2024a; b); Wu et al. (2024b). In each communication round $r\in[R]$ with the global model $\phi_{r}$, the following steps are conducted:
@@ 4 FedBis: A Vanilla and Feasible Framework for Achieving Federated RLHF
The proposed FedBis provides a simple yet effective solution through a three-step process to enable model fine-tuning with clients’ preference data while respecting clients’ privacy, as depicted in Figure 2: In the first step, each client builds their own preference dataset, which is originated from the daily queries (a.k.a. prompts or instructions) to a pretrained language model (LM), the model generated a pair of responses to each query, and the client chooses the preferred one. After the construction of the preference dataset, each client independently trains the selector, and the server aggregates the selectors into a global one and broadcasts it to the clients. This communication process repeats for a total of $R$ times, and the clients keep the preference dataset unchanged during the training. Afterward, we utilize the well-trained binary selector to enhance the performance of LLM. Specifically, we assume the server holds a set of instructions, together with pairwise responses generated by an LLM. Then, we build a preference dataset with the help of the binary selector and boost the LLM by means of DPO Rafailov et al. (2023). The rest highlights the key steps of these two phases, while Appendix A offers a detailed description of the proposed FedBis.
@@
While the proposed FedBis effectively achieves RLHF in FL with low computational costs, there are two key limitations that remain unsolved. The first limitation is preference heterogeneity. The proposed work effectively aggregates client preferences via clients’ preference models, but it does not address the preference gap across clients when there is a huge difference in client preferences and prompts. As a result, the clients optimize their local models for their own data instead of a global objective. This leads to global aggregation diverging from the global optimum, which lengthens the training time to obtain the desired performance Karimireddy et al. (2020); Wu et al. (2023a; b). Another limitation is reward hacking. The proposed FedBis relies on training a single selector and using it to fine-tune the LLM. This procedure introduces an adversarial dynamic where the model "cheats" the selector into favoring certain responses without genuinely improving. Eventually, as the LLM is trained across more iterations, its performance degrades significantly, making the approach inefficient and unsustainable. In the coming section, we propose a new algorithm that is able to address both limitations while maintaining low computational costs.
@@ 5.1 Algorithm Design
The proposed FedBiscuit consists of two phases: 1) We train each selector for a couple of rounds so that all $U$ selectors have basic capacities in selecting the preferred completion, and 2) we divide the clients into disjoint clusters of size $U$ and train each binary selector with a specific cluster.
@@
In the beginning, we initialize each binary selector $\phi_{u}(u\in[U])$ with an identical pretrained LLM $\tilde{\phi}$. Subsequently, starting from $u=0$, we train a selector $\phi_{u}$ for $T$ consecutive communication rounds following the steps of FedBis: In each communication round, the server samples a subset of client $\mathcal{A}$ and broadcasts the selector $\phi_{u}$ to them. Each client $m\in\mathcal{A}$ then locally trains the selector for $K$ iterations using the dataset $\mathcal{D}_{m,train}$. At the end of the communication round, the server aggregates and updates the selector $\phi_{u}$ via Equation (8). After completing the training of $\phi_{u}$, the server initiates the training of the next selector $\phi_{u+1}$ by repeating the above steps until all selectors are trained.
@@ 6 Federated Human Preference Benchmark
We reconstruct the public dataset SHP, which comprises numerous questions from Reddit posts and their corresponding user answers Ethayarajh et al. (2022). The preference indicator is based on the number of likes an answer receives. Following the training of StreamSHP Ethayarajh et al. (2022), we utilize the data with no more than 512 tokens. Given that the dataset spans 18 domains, we partition the dataset using a Dirichlet distribution with a parameter of 0.3, ensuring that no questions overlap between clients. In our experiment, we consider training the binary selector with 200 clients, which is a common setting when evaluating the performance of an FL algorithm Jhunjhunwala et al. (2023). Figure 3 visualizes the data distribution on the selected clients. For the RLHF process, we incorporate 2.6K Reddit questions and 44.6K SafeRLHF prompts Dai et al. (2023).
@@ 7.3 Quantitative Evaluation on QA Task
Table 2: Performance under QA task using AlpacaEval 2.0. “LC” in the table means “length-control.” FedBis and FedBiscuit adopt the reward model of Gemma-2B fine-tuned on SHP dataset. Bold highlights the best result under each column, while Underline visualizes the best result under different sources of unlabeled prompts.
@@ 8 Conclusion
In this work, we explore a feasible framework to employ federated learning technique to enable large-scale user preference collection for RLHF without jeopardizing user privacy. Specifically, we train a binary selector across different clients using their local preference datasets, and then use the well-trained selector to align an LLM with human preferences. We propose two approaches to enable selector training: FedBis and FedBiscuit. FedBis provides a framework to train a single selector, while FedBiscuit ensembles multiple selectors to address preference heterogeneity and reward hacking. We conduct empirical studies with the proposed federated human preference datasets to validate our statements and demonstrate the superiority of FedBis and FedBiscuit when aligning Gemma-2B and LLaMA-2-7B with human preference.
Ethics Statement
@@ A.1 Preference Modeling
We consider a practical and efficient FL scenario where not all clients but only a sampled subsets of clients participate in each communication round Yang et al. (2020). Before the commencement of FL training, we initialize the binary selector with a pretrained LLM such as LLaMA-2 Touvron et al. (2023), and set the hyperparameters. An FL algorithm requires multiple communication rounds and consists of three phases in each round, i.e., model broadcast, local training, and global aggregation. Following this paradigm, we design FedBis and optimize the selector $\phi$, i.e., in the communication round $r\in[R]$, as discussed as follows.
@@
This aggregation method, based on Li et al. (2019) where the clients are uniformly sampled to train a global model, ensures consistency with Problem (1) in mathematical expectation.

### v3

@@ 1 Introduction
Large language models (LLMs) exhibit broad knowledge coverage as they are pretrained on a large corpus Zhao et al. (2023a); Singhal et al. (2023); Wang et al. (2023a); Liu et al. (2024a). To elevate the quality of their generated content to match the professionalism and readability of human writing, a common approach is to fine-tune these models using reinforcement learning with human feedback (RLHF) Ziegler et al. (2019); Christiano et al. (2017); Ouyang et al. (2022). This process relies on preference datasets, which are constructed through two main methods: human evaluation Bai et al. (2022); Ganguli et al. (2022); Stiennon et al. (2020) and ChatGPT-based ranking Dubois et al. (2024b). In the human-effort approach, a set of instructions (a.k.a. prompts or user queries) is paired with multiple model completions (a.k.a. generated responses), and a team of labelers ranks completions of each instruction from best to worst. In the ChatGPT-based method, the dataset is developed by feeding ChatGPT with multiple inputs encompassing an instruction and a pair of completions, where it selects the superior completion for each input.
@@
As LLMs are deployed to serve diverse users, there could be a gap between the preferences of real-world users and those of lablers/ChatGPT, hindering the LLM’s ability to generate responses that align with the users’ preferences. Therefore, there is a need for a preference dataset that accurately reflects real world users’ preferences in order to enhance the ability of LLM in content generation.
@@
To address privacy concerns, we propose utilizing Federated Learning (FL) techniques to enable large-scale preference collection from diverse real-world users without requiring them to transmit their preference data to a central server. In our design, we adopt FedAvg, a well-established FL algorithm Konečn‘y et al. (2016); McMahan et al. (2017), to learn user preferences without directly collecting their data. In line with the reward model training used in RLHF Stiennon et al. (2020); Ouyang et al. (2022); Rafailov et al. (2023), each LLM user works as a FL client and locally trains a reward model that can evaluate the quality of model completions. The server then aggregates these client-trained reward models into a global model. While this approach appears effective, we observe the following two limitations during training:
  • Excessive Computation Overhead: Typically, the reward model is a regression model designed to output a scalar value that represents the quality of a model completion Stiennon et al. (2020). Its optimization is based on comparing reward differences between preferred and dispreferred completions, where preferred completions should receive higher rewards. However, during optimization, each data sample requires the retention of two computation graphs (one for the preferred and one for the dispreferred completion) throughout the forward and backward passes. This results in significant computational overhead and heavy GPU demands.
@@ 2 Related Work
Recent studies have increasingly focused on fine-tuning large language models (LLMs) using federated datasets Sun et al. (2024); Zhang et al. (2024; 2023); Yi et al. (2023); Qin et al. (2023; 2024); Bai et al. (2024). However, these approaches often suffer from high computation and communication costs due to the necessity of training and synchronizing the model with clients. To mitigate these issues, lightweight methods such as black-box fine-tuning Sun et al. (2023); Lin et al. (2023) and offsite-tuning Wu et al. (2024a); Kuang et al. (2023) have emerged. Despite their advancements, these methods primarily focus on fine-tuning LLMs for specific downstream tasks, neglecting user preferences in the generated responses. A recent benchmark, OpenFedLLM Ye et al. (2024), introduces FedDPO, which allows federated clients to optimize their local LLMs using DPO loss. While this approach can potentially align LLMs with human preferences, it faces three key challenges: excessive computational overhead, preference heterogeneity, and the risk of reward hacking. To address these limitations, our work aims to enable LLMs alignment with a feasible and sustainable training framework in FL.
@@
RLHF typically involves supervised fine-tuning, reward modeling, and reward optimization, initially proposed by Christiano et al. (2017). Proximal Policy Optimization (PPO) Schulman et al. (2017) is a common RLHF algorithm, yet it struggles with instability, inefficiency, and high resource demands Choshen et al. (2019); Engstrom et al. (2020). These challenges have led to the development of alternative methods, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and others Dong et al. (2023); Zhao et al. (2023b); Azar et al. (2024); Ethayarajh et al. (2024); Gulcehre et al. (2023), which offer more stable and efficient solutions. However, these methods typically operate within a centralized training framework, where the LLM owner retains control over the preference data. In contrast, our work aims to expand data sources and integrate real user preferences without directly collecting their personal data.
@@ 3.1 Federated Learning (FL)
FL is a distributed training paradigm where a server coordinates various clients toward the same goal, i.e., training a generalized model $\phi\in\mathbb{R}^{d}$ Konečn‘y et al. (2016); McMahan et al. (2017). Consider an FL system with $M$ clients. Denote the weight of client $m$ as $p_{m}$ such that $\sum_{m\in[M]}p_{m}=1$, and we aim to optimize the following objectives:
$\min_{\phi\in\mathbb{R}^{d}}\quad F(\mathbf{\phi})\overset{\triangle}{=}\sum_{m\in[M]}p_{m}F_{m}(\mathbf{\phi}),$ (1)
where $F_{m}(\phi)$ is the expected loss on client $m$ given the model $\phi$. As a classical FL algorithm, FedAvg can solve the optimization problem by multiple communications between the server and the clients, followed by a number of recent works Wang et al. (2021; 2022; 2023c; 2024b; 2024c); Zhang et al. (2021). In each communication round $r\in[R]$ with the global model $\phi_{r}$, the following steps are conducted:
@@ 4 FedBis: A Vanilla and Feasible Framework for Achieving Federated RLHF
The proposed FedBis provides a simple yet effective solution through a three-step process to enable model fine-tuning with clients’ preference data while preserving clients’ privacy, as depicted in Figure 2: In the first step, each client builds their own preference dataset, which is originated from the daily queries (a.k.a. prompts or instructions) to a pretrained language model (LM), the model generated a pair of responses to each query, and the client chooses the preferred one. After the construction of the preference dataset, each client independently trains a binary selector, and the server aggregates these selectors into a global one and broadcasts it to the clients. This communication process repeats for a total of $R$ times, and the clients keep the preference dataset unchanged during the training. In contrast to traditional RLHF methods that train a regression-based reward model, this step employs a binary selector to alleviate computational constraints on edge devices Wu et al. (2023b; 2024b); Wang et al. (2023b; 2024a); Li et al. (2024). Additionally, the selector enables comprehensive pairwise comparisons between completions, thereby facilitating more efficient training Liu et al. (2024b); Dubois et al. (2024b). Afterward, we utilize the well-trained binary selector to enhance the performance of LLM. Specifically, we assume the server holds a set of instructions, together with pairwise responses generated by an LLM. Then, we build a preference dataset with the help of the binary selector and boost the LLM by means of DPO Rafailov et al. (2023). This section highlights the key steps of the proposed FedBis, while Appendix A offers a detailed description.
@@
While the proposed FedBis effectively achieves RLHF in FL with low computational costs, there are two key limitations that remain unsolved. The first limitation is preference heterogeneity. The proposed work effectively aggregates client preferences via clients’ preference models, but it does not address the preference gap across clients when there is a huge difference in client preferences and prompts. As a result, the clients optimize their local models for their own data instead of a global objective. This leads to global aggregation diverging from the global optimum, which lengthens the training time to obtain the desired performance Karimireddy et al. (2020); Wu et al. (2023a). Another limitation is reward hacking. The proposed FedBis relies on training a single selector and using it to fine-tune the LLM. This procedure introduces an adversarial dynamic where the model "cheats" the selector into favoring certain responses without genuinely improving. Eventually, as the LLM is trained across more iterations, its performance degrades significantly, making the approach inefficient and unsustainable. In the coming section, we propose a new algorithm that is able to address both limitations while maintaining low computational costs.
@@ 5.1 Algorithm Design
The proposed FedBiscuit consists of two phases: 1) We train each selector for a couple of rounds so that all $U$ selectors have basic capacities in selecting the preferred completion, and 2) we divide the clients into disjoint clusters of size $U$ and train each binary selector with a specific cluster. A full pseudocode implementation is provided in Algorithm 1.
@@
In the beginning, we initialize each binary selector $\phi_{u}(u\in[U])$ with an identical pretrained LLM $\tilde{\phi}$. Subsequently, starting from $u=0$, we train each selector $\phi_{u}$ for $T$ consecutive communication rounds following the steps of FedBis: In each round, the server samples a subset of clients $\mathcal{A}$ and broadcasts the selector $\phi_{u}$ to them. Each client $m\in\mathcal{A}$ then locally trains the selector for $K$ iterations using the dataset $\mathcal{D}_{m,train}$. At the end of the communication round, the server aggregates and updates the selector $\phi_{u}$. After completing the training of $\phi_{u}$, the server initiates the training of a new selector $\phi_{u+1}$ by repeating the above steps until all selectors are trained.
@@ 6 Federated Human Preference Benchmark
We reconstruct the public dataset SHP, which comprises numerous questions from Reddit posts and their corresponding user answers. The preference indicator is based on the number of likes an answer receives. Following the training of StreamSHP Ethayarajh et al. (2022), we utilize the data with no more than 512 tokens. Given that the dataset spans 18 domains, we partition the dataset using a Dirichlet distribution with a parameter of 0.3, ensuring that no questions overlap between clients. In our experiment, we consider training the binary selector with 200 clients, which is a common setting when evaluating the performance of an FL algorithm Jhunjhunwala et al. (2023). Figure 3 visualizes the data distribution on the selected clients. For the RLHF process, we incorporate 2.6K Reddit questions and 44.6K SafeRLHF prompts Dai et al. (2023).
@@ 7.3 Quantitative Evaluation on QA Task
Table 2: Performance under QA task using AlpacaEval 2.0. "LC" in the table means "length-control." FedBis and FedBiscuit adopt the reward model of Gemma-2B fine-tuned on SHP dataset. Bold highlights the best result under each column, while Underline visualizes the best result under different sources of unlabeled prompts.
@@ 8 Conclusion
In this work, we explore a feasible framework to employ a federated learning technique to enable large-scale user preference collection for RLHF without jeopardizing user privacy. Specifically, we train a binary selector across different clients using their local preference datasets, and then use the well-trained selector to align an LLM with human preferences. We propose two approaches to enable selector training: FedBis and FedBiscuit. FedBis provides a framework to train a single selector, while FedBiscuit ensembles multiple selectors to address preference heterogeneity and reward hacking. We conduct empirical studies with the proposed federated human preference datasets to validate our statements and demonstrate the superiority of FedBis and FedBiscuit when aligning Gemma-2B and LLaMA-2-7B with human preference.
Ethics Statements
@@
Acknowledgements
The authors would like to thank the anonymous reviewers for their constructive comments. This work is supported in part by the US National Science Foundation under grants NSF CNS-2154059, NSF IIS-2141037, and NSF IIS-2226108. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 Preference Modeling
We consider a practical and efficient FL scenario where not all clients but only a sampled subsets of clients participate in each communication round Yang et al. (2020); Wu et al. (2023a); He et al. (2023); Fang et al. (2024). Before the commencement of FL training, we initialize the binary selector with a pretrained LLM such as LLaMA-2 Touvron et al. (2023), and set the hyperparameters. An FL algorithm requires multiple communication rounds and consists of three phases in each round, i.e., model broadcast, local training, and global aggregation. Following this paradigm, we design FedBis and optimize the selector $\phi$, i.e., in the communication round $r\in[R]$, as discussed as follows.
@@
This aggregation method, based on Li et al. (2019) where the clients are uniformly sampled to train a global model, ensures consistency with Problem (1) in mathematical expectation. To prevent the server from inferring a client’s preferences from their local model, a practical solution is secure aggregation, which ensures the server only accesses the aggregated results Bonawitz et al. (2016; 2017); Kairouz et al. (2021); Chen et al. (2022); So et al. (2022; 2023); Rathee et al. (2023).